OpenStudy (anonymous):

A forest ranger spots a fire from a 38-foot tower. The angle of depression from the tower to the fire is 11 degrees. To the nearest foot, how far is the fire from the base of the tower? Show the steps you use to find the solution.

OpenStudy (anonymous):

 tan(38 degrees) = 38/F ---> \(\large F=\frac{38}{tan(38)} \) can you take it from here?